OpenStudy (anonymous):

Greetings, I would like to know what the actual infinitisimal magnetic flux (dΦ) of a conductive ring is. More details below.

OpenStudy (anonymous):

* I ask for the infinitesimal flux of the conductive ring, and after that the EMF * B =Cxxˆ−Cyyˆ * V is constant on the x axis (we have only X axis velocity) * (0,0) center of the ring My results: for dφ I found this to be like Cx2πadx and EMF -2πa(v^2)t

OpenStudy (anonymous):

* note that dΦ = \[\vec{B} \vec{dS}\] or \[\int\limits_{strip}^{}\vec{B} d\vec{A}\] where strip is the orthogon strip that connects the ring in thr two phases (before and after)

OpenStudy (vincent-lyon.fr):

Is axis of the ring parallel to x-axis?

OpenStudy (anonymous):

the axis of the ring is the x-axis - (0,0) centered ring

OpenStudy (anonymous):



OpenStudy (anonymous):

is the normal area vector ring perpendicular to B or parallel??

OpenStudy (anonymous):

if perpendicular then 0 else it will depend on the order of integration..

OpenStudy (anonymous):

wait check the B field above, it is neither perpendicular nor parallel

OpenStudy (anonymous):

is there one or two rings??

OpenStudy (anonymous):

one ring during an infitesimal motion at x-axis, let's say this to be dx

OpenStudy (anonymous):

x axis is its centroidal axis?? sorry cant see where is the disk moving..

OpenStudy (anonymous):

which products a dΦ οf \[\vec{B} \vec{dS}\] or \[\int\limits_{strip}^{}\vec{B}\vec{dA}\]

OpenStudy (anonymous):

centroidal

OpenStudy (anonymous):

**where strip is the infitesimal orthogonal strip with sides dl=dx and dl=adθ don't forget that during infitesimal motions we ignore the Φ of the main surfaces because thry are the same and can't be included in dΦ so we have only to do with side flux

OpenStudy (anonymous):

(**i opened this topic just for discussion and to check my answers to this problem)

OpenStudy (anonymous):

i cant see what the fig is like..but project the ring on XY plane and then apply the formula by taking a stripparallel to x axis..

OpenStudy (anonymous):

just imagine 2 rings on almost the same location (distance=dx) and that "dx line" connecting them.. then imagine the surface they create (a cylinder surface)

OpenStudy (anonymous):

also a second way to view this according to my opinion is to consider a small part of this "cylinder" which makes a ring again. Then create very small rectangular parts on it's surface with side 1 = dx and side 2 = dl (but dl is a polar motion due to the electric current which advances on the ring) so side 2 = dl = adθ

OpenStudy (anonymous):

if anyone disagrees just let me know

OpenStudy (vincent-lyon.fr):

Magnetic flux through ring is \(\Phi=\pi a^2Cx\) and electromotive force is \(e=-\pi a^2Cv\) where \(a\) is the radius of the ring. I do not find wise to calculate the magnetic flux 'cut' by the ring. you will need to integrate \(\cos ^2\theta \:d\theta\)

OpenStudy (anonymous):

yes but we need dΦ not Φ, and it comes up to a linear flux. Check Griffiths Introduction to Electrodynamics of Faraday's Law/Magnetic Flux part, they say that we can work like this to find the dΦ

OpenStudy (anonymous):

S (main) surface is stable so the dΦ comes from linear flux which comes from infinitesimal rectangulars

OpenStudy (vincent-lyon.fr):

Of course you can, but it is much more complicated, than working the actual flux through the circuit. You will get a similar expression: \(d\Phi=-\pi a^2C dx\)

OpenStudy (anonymous):

Thank you for for reply Vincent. According to Griffith's Introduction to Electrodynamics, the dΦ comes from the following integral: Φstripe where stripe is the "line" that sticks the two phases together. Imagine something like the following (IMPORTANT: the field is not homogeneous)  this shape represents the magnetic field and the infimitesial motion of the ring (**continue bellow)